OpenStudy (anonymous):

. Find the x and y intercepts. Please show all of your work. -3x + 8y = 24

OpenStudy (anonymous):

yint=3

OpenStudy (anonymous):

x=-8

OpenStudy (anonymous):

3x+24=3y x+8=y

OpenStudy (anonymous):

y intercept is 8 and x intercept is -8

OpenStudy (anonymous):

x intercept is when y =0 y intercept is when x=0

OpenStudy (anonymous):

Yeah... except the question asks for 8y NOT 3y so the y intercept will be 3 but yes the x intercept is -8 and yes you find the intercepts by setting the opposite variable to 0.

OpenStudy (anonymous):

so it would be this (-8,0)(0,-8)

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

ill plot it for you

OpenStudy (anonymous):

Oh u are right I am sorry for confusing u

OpenStudy (anonymous):

simple one equation so just plug in the value for x = 0 once and y = 0 other time u will get the x and y intercepts !!!!!

OpenStudy (anonymous):

3x+24=8y 3/8x+3=y (3/8)(0)+3=y y=3 y intercept 3/8x+3=0 3/8x=-3 3x=-24 x=-8 x interecept

OpenStudy (anonymous):

now that I understand.. thanks

OpenStudy (anonymous):

when x = 0 -3x + 8y = 24 => -3*0 + 8y = 24 => y = 3 when y = 0 -3x + 8*0 = 24 => x = -8 intercepts are (0, 3) and (-8,0)

OpenStudy (anonymous):

I always give everyone the wrong answer and confuse them!

OpenStudy (anonymous):

haahahha u dont feel like that..at least u dont use softwares for solving